Homework: Section 2.2 Score: 0 of 1 pt 2.2.27 Homework Save 5 of 11 (1 complete) HW Score: 9 09%, 1 of 11 pts Question Help between two health clubs Club A offers membership for a fee of $23 plus a monthiy fee of $18 Club B offers membership for a fee of $21 plus a monthly fee of $19 After how many months will the total cost of each health club be the same? What will be the total cost for each club? in months the total cost of each health club will be the same OM
Answer #1

for club-A.......23 plus 18 per month......23+18x...........where x is number of months

for club-B.......21 plus 19 per month......21+19x...........where x is number of months

when both are equal....

\large 23+18x=21+19x

\large 19x-18x=23-21.

\large {\color{Red} x=2}

after 2 months total cost will be same

and total cost is

\large 23+18x

\large =23+18(2)

\large =23+36

\large {\color{Red} =59}